OpenStudy (anonymous):

Find the x- and y-intercepts for the following equation. y = 2x + 4 x-intercept (x, y) = (, ) y-intercept (x, y) = (, )

OpenStudy (therealmeeeee):

Simplifying 4x + -2y + -6 Reorder the terms: -6 + 4x + -2y Factor out the Greatest Common Factor (GCF), '2'. 2(-3 + 2x + -1y) Final result: 2(-3 + 2x + -1y) set equal to zero Simplifying -3 + 2x + -1y = 0 Solving -3 + 2x + -1y = 0 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '3' to each side of the equation. -3 + 2x + 3 + -1y = 0 + 3 Reorder the terms: -3 + 3 + 2x + -1y = 0 + 3 Combine like terms: -3 + 3 = 0 0 + 2x + -1y = 0 + 3 2x + -1y = 0 + 3 Combine like terms: 0 + 3 = 3 2x + -1y = 3 Add 'y' to each side of the equation. 2x + -1y + y = 3 + y Combine like terms: -1y + y = 0 2x + 0 = 3 + y 2x = 3 + y Divide each side by '2'. x = 1.5 + 0.5y Simplifying x = 1.5 + 0.5y

OpenStudy (anonymous):

i needed to find the x & y intercepts @TheRealMeeeee

OpenStudy (therealmeeeee):

thats it

OpenStudy (anonymous):

is it xintercept 1,5 and y intercept 0,5 ?

OpenStudy (anonymous):

@TheRealMeeeee